OpenStudy (anonymous):

You wonder how many cars there are in your area. You call up the Ford dealership and ask how many Fords they’ve sold. They report that they’ve sold a total of 15,000 cars. You then take a lawn chair and camp out on a busy road and count cars. You count 1200 cars and note that 700 of them are Fords. Estimate the number of cars in your area.

OpenStudy (mimi_x3):

@dan815 youve got experience in this field

OpenStudy (anonymous):

Do you have an answer for this?

OpenStudy (dan815):

lol

OpenStudy (dan815):

700/1200 are ford so

OpenStudy (dan815):

15000/x = 700/1200

OpenStudy (dan815):

7*(2+1/7)=15 so

OpenStudy (dan815):

and 12 is 7+5 = 7(1+5/7)

OpenStudy (dan815):